HETAV PATEL

Grade 11,

If the electric flux entering and leaving an enclosed surface respectively is ? 1 and ? 2 , the electric charge inside the surface will be (a) (? 2­ – ? 1 )e 0 (b) (? 1­ – ? 2 )/e 0 (c) (? 2­ – ? 1 )/e 0 (d) (? 1 – ? 2 )e 0

If the electric flux entering and leaving an enclosed surface respectively is ?1 and ?2, the electric charge inside the surface will be
 
(a) (? – ?1)e0
(b) (? – ?2)/e0
(c) (? – ?1)/e0
(d) (?1 – ?2)e0

Grade:6

2 Answers

bharat bajaj IIT Delhi
9 years ago
Electric flux = q/eo
?1 - ?2 = q/eo
Electric charge = (?1-?2)eo
D) is the answer
THanks
Bharat Bajaj
IIT Delhi
askiitians faculty
Shivani Yadav
6 years ago
( Phi1-Phi2)eoPhi2-phi1= qenc/eoqenc=(phi1-phi2)eoPhi2-phi1 is equal to qenc /eoSo,qenc will be equal to phi2-phi1 eo